One of the uses of titanium is on the structure of the bodies of planes since it is strong despite of being lightweight. Titanium has a density of 4.55 g/cm3. What is the diameter (in millimeters) of a cylinder of titanium that is 7.75 cm long with a mass of 0.153kg
1
Expert's answer
2020-11-17T10:38:42-0500

p=m/V.

V=m/p.

V= hπr^2.

r=√(V/hπ)

d=2r.

V=153/4.55 = 33,6263736 cm3.

r=√(33,6263736/7.75 x 3.14)= 1,17550334 cm.

d= 1,17550334 x 2 = 2,35100668 cm.